Suppose Ford sold an issue of bonds with a 16-year maturity, a $1200 par value, a 10% coupon rate, and semiannual interest paplease do it asap
Answer #1

Answer A)

Valucof Bond = Coupon + – 1- (1+r) Maturity Value (1+r)

where r is the the rate of return i.e. 7%/ 2 = 3.5 % _ _ _ (Semi-Annual)

n is the number of compounding period i.e. (16-2) * 28 periods

Coupon is interest (semi Annual) = 10%/2 * 1200 = $ 60

1- Valueof Bond = 60* — (1+0.035) 25 0.035 + 1200 (1 + 0.035 28

= 60* 17.667019 + 1200* 0.3816543

= $ 1518

Answer B)

Valucof Bond = Coupon + – 1- (1+r) Maturity Value (1+r)

where r is the the rate of return i.e. 14%/ 2 = 7% _ _ _ (Semi-Annual)

n is the number of compounding period i.e. (16-2) * 28 periods

Coupon is interest (semi Annual) = 10%/2 * 1200 = $ 60

1- Valueof Bond = 60* — (1+0.07) 0.07 1200 (1 + 0.07 28

= 60* 12.1371 + 1200* 0.150402

= $ 908.71

Answer C)

Current Yield = Coupon / Bond Price

= 120/783.58

= 15.31%

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.